8. Which of the following is correct? OSMC = PMC - MD OPMB = SMB + EMB OSMB = SMC + PMB OSMC = PMC + MD
Answer #1

Social marginal cost is the sum of private marginal cost and marginal damage.

SMC = PMC + MD.

Social marginal benefit is the sum of private marginal benefit and external marginal benefit.

SMB = PMB + EMB

=> PMB = SMB - EMB.

Answer: Option (D) i.e., SMC = PMC + MD